# Multi-Robot Coordination Induced in Hazardous Environments through an Adversarial Graph-Traversal Game

### v1

@@ A More Practical Example
Here, we discuss how these strategies we’ve discussed on smaller graphs fit into a larger graph. Figure 6 shows portions of an example trajectory under equilibrium policies for a larger graph of $N=10$ nodes, with $M=4$ blue robots, and a maximum ammo count for the red player of $\overline{\alpha}=5$, while Table I contains more details. Both players employ mixed strategies, so this is only one possible outcome for these initial conditions. For example, at $t=1$, the blue player’s policy can be described by $[\text{Pr}(p^{+}=[1,1,5,6]^{T}),\text{Pr}(p^{+}=[1,1,6,6]^{T}),\text{Pr}(p^{+}=[1,2,6,6]^{T})]=[0.4928,0.3482,0.1590]$ and the red player’s is $\pi_{\text{red}}^{*}(s(1))=[0.6536,0.0818,0.2646]^{T}$. Additionally, the blue player makes use of some self-loops to allow some of its robots to wait, and the blue robots also split up. The value per robot is $\tilde{V}_{4}(S(0))=14.4667$, while the value per robot for a single robot is higher at $\tilde{V}_{1}(S(0))=15.2474$, again suggesting that having more robots is helpful for the blue player. Finally, this value is between the lower and upper bounds for the value, $\underline{V}(S(0))=12$ and $\overline{V}(S(0)=80$, as we expect.
@@ 6 $[10,10,10,10]^{T}$ ${G^{3}}$ $3$ 62
TABLE I: Example Nash equilibrium trajectory for the graphs shown in Figure 6
@@ V-B Statistical Results
To statistically demonstrate some properties of this game and its solution, we constructed graphs in a random fashion. More specifically, we used a modification of the canonical Erdős–Rényi model [20]. For a given set of nodes $\mathcal{V}$ of size $N_{\text{max}}$, the probability that any given pair of nodes $(i,j)\in\mathcal{V}\times\mathcal{V}$, for $i\neq j$, is in the edge set of the graph is a constant $\varphi\in[0,1]$. We also consider directed edges and possible self-loops, and we take $\varphi=0.5$. To ensure the goal can be reached from all nodes, we find the longest directed distance between any two nodes, and assign one of those ending nodes to be the goal node and a corresponding starting node to be node $1$. Then, we remove any nodes, after the initial generation, from which the goal node cannot be reached, which results in a graph with $N\leq N_{\text{max}}$ nodes. Finally, we add a self-loop to the goal node $N$, if necessary. We set $K=3$, and randomly assign edge weights for each graph ${G^{k}}$ so that, for $(i,j)\in\mathcal{E}$ and $i\neq j$, $(W^{1}_{ij},W^{2}_{ij},W^{3}_{ij})$ is a permutation of the set $\{2,4,8\}$, and each permutation is chosen with equal probability. The self-loop weights are assigned according to $W^{k}_{ii}=1$ for $(i,i)\in\mathcal{E}$ and $i\neq N$ and $W^{k}_{NN}=0$. Using this method, $100$ graphs were generated for each initial graph size $N_{\text{max}}$ between 4 nodes and 8 nodes.
@@
We consider a common initial condition, where all blue robots are on node 1 in graph 1, so that $p(0)=\mathbf{1}_{M}$ and ${G}(0)={G^{1}}$, and the red player has ammo $\alpha(0)=6$. In Figure 7, the values for this initial condition have been normalized so that the lower bound $\underline{J}\left(s(0)\right)$ is zero and the upper bound $\overline{J}\left(s(0)\right)$ is one in the plot. Some trivial graphs have been excluded from the results because $\underline{J}\left(s(0)\right)=\overline{J}\left(s(0)\right)$ on them. The bounds are respected in all cases, and are sometimes tight. Additionally, the median and mean normalized value exhibits a downward trend as the number of nodes increases. This is likely because more nodes makes it more probable that the blue player will have more options for paths to take to the goal, allowing it to gain an advantage through mixed policies, for example.
@@
Figure 7: This shows the value per robot $\tilde{V}_{M}(S(0))$, from a common initial condition, over about $500$ random graphs, which has been normalized so that the lower bound is zero and the upper bound is $1$. For each different number of blue robots, the value is plotted against the number of nodes in the random graphs. The plus sign indicates the mean, the center horizontal line indicates the median value, the middle $50\%$ of the values are inside the box, the whiskers indicate the furthest non-outlier values, and the circles indicate the outliers. The values are always between the upper and lower bounds, there is a general downward trend in the median and mean value as the number of nodes increases, and the mean decreases with increasing numbers of robots.
Effect of the Number of Blue Robots
@@
As discussed in the preceding sections, we expect that having more blue robots will decrease the cost per robot. Figure 7 also plots the values per robot $\tilde{V}_{M}(S(0))$ for the random graphs for different numbers of robots. For the sake of computation time, the game was solved for $M=4$ only for graphs with $N\leq 7$. We observe that the mean value per robot does decrease as the blue player’s number of robots increases for these graphs, for each graph size. This is evidence that having more robots is indeed an advantage for the blue player.
Effect of the Red Player’s Ammo Count
@@
Figure 8 shows the effect of the red player’s ammo count on the normalized game value. Here, the same initial position for the blue player is considered, but the initial amount of ammo for the red player is varied. The average of the normalized values is taken over all the random graphs and plotted against the ammo count. As expected, the value is nondecreasing with respect to the ammo count, and the value plateaus around 3 or 4 ammo, below the upper bound for the value. This suggests that having more ammo stops benefiting the red player in terms of the value, once it is high enough.
@@
Figure 8: This shows the average, over about $500$ random graphs, cost from the starting node to the ending node, as a function of ammo. It’s been normalized so that the lower bound is zero and the upper bound is $1$, but the plot bounds are only between $0.6$ and $0.8$ to save space. The value appears to be a nondecreasing function of the ammo and it asymptotically approaches a value which is significantly lower than the upper bound.
@@ VI Conclusions
To handle uncertainty in a time-varying environment, we have formulated a novel zero-sum game where the blue player’s robots move through a graph with time-varying costs to reach a goal, while the red player controls those costs. This can be appropriate for robots moving through hazardous environments, either to model the worst-case scenario or to handle an actual adversary. Due to the complexity of the game, we provided security strategies for both players, and demonstrated the game solution through numerical methods, with a theoretical guarantee that the blue robots reach the goal. We provided methods to simplify computation, and find that both players benefit from mixed strategies, while the blue player benefits from having multiple robots which use coordinated strategies. These include splitting up and waiting to traverse risky edges synchronously, ensuring at least some take the cheaper paths. In future work, we may analytically solve the game in some specific cases. Additionally, this work has made the strong assumption of perfection information for the blue player. Future work can include formulating and solving games of incomplete information, where the blue player is uncertain of the red player’s capabilities, and imperfection information, where the blue robots may only be aware of changes made to the graph close to them. This would help model the more realistic case of limited perception for the robots.

### v2

@@ A More Practical Example
Here, we discuss how these behaviors fit into a larger graph. Figure 6 shows portions of an example trajectory under equilibrium policies for a larger graph of $N=10$ nodes, with $M=4$ blue robots, and a maximum ammo count for the red player of $\overline{\alpha}=5$, while Table I contains more details. Both players employ mixed strategies, so this is only one of the possible outcomes. For example, at $t=1$, the blue player mixes so that its next possible positions are $[1,1,5,6]^{T}$, $[1,1,6,6]^{T}$, and $[1,2,6,6]^{T}$, so that some robots wait and some robots may split up, while the red player’s policy is mixed between choosing the each of the three graphs. The value per robot is $\tilde{V}_{4}(S(0))=14.4667$, while it is higher for a single robot at $\tilde{V}_{1}(S(0))=15.2474$, again suggesting that having more robots is helpful for the blue player. Finally, this value is between the lower and upper bounds for the value, $\underline{J}\left(S(0)\right)=12$ and $\overline{J}\left(S(0)\right)=80$, as we expect.
@@ 6 $[10,10,10,10]^{T}$ ${G^{3}}$ $3$ 62
TABLE I: Example NE trajectory for Figure 6
@@ V-B Statistical Results
To statistically demonstrate some properties of this game and its solution, we constructed graphs using a modification of the canonical Erdős–Rényi model [27]. For a set of nodes $\mathcal{V}$ of size $N_{\text{max}}$, the probability that any pair of nodes $(i,j)\in\mathcal{V}\times\mathcal{V}$ is in the directed edge set of the graph is a chosen constant $0.5$. Node $1$ and the goal node are the nodes with the longest (unweighted) distance between them, and any nodes from which the goal node cannot be reached are pruned, so the graph size is $N\leq N_{\text{max}}$. Finally, a self-loop is added to the goal node $N$. We set $K=3$, and randomly assign edge weights for each graph ${G^{k}}$ so that, for $(i,j)\in\mathcal{E}$ and $i\neq j$, $(W^{1}_{ij},W^{2}_{ij},W^{3}_{ij})$ is a permutation of the set $\{2,4,8\}$, chosen with uniform probability. The self-loop weights are all $1$ except for the goal node’s. Using this method, $100$ graphs were generated for each initial graph size $N_{\text{max}}$ between 4 nodes and 8 nodes.
@@
The common initial condition is $p(0)=\mathbf{1}_{M}$ and ${G}(0)={G^{1}}$, and the red player has ammo $\alpha(0)=6$. In Figure 7, the values for this initial condition, for 5, 6, and 7 node graphs, have been normalized so that the lower bound $\underline{J}\left(s(0)\right)$ is 00 and the upper bound $\overline{J}\left(s(0)\right)$ is $1$ in the plot. Trivial graphs where $\underline{J}\left(s(0)\right)=\overline{J}\left(s(0)\right)$ have been excluded from the results. Additionally, the approximate expected cost is shown for alternative blue strategies, which are the security strategy defined in (III-C1), and a naive strategy where robots follow the shortest path on a “best case” graph, where all edges have the smallest cost. These have the same expected cost for any number of robots. The bounds are respected and are sometimes tight for all NE policies and the security strategy, while the naive strategy sometimes achieves much higher normalized costs for the blue player. Compared to the baseline of the security strategy, the blue NE policy has lower costs, and both outperform the naive blue strategy. Additionally, the mean normalized value exhibits a downward trend as the number of nodes increases. More nodes may give the blue player more options to reach the goal, allowing it to gain an advantage through mixed policies.
@@
Figure 7: This shows the value per robot $\tilde{V}_{M}(S(0))$ over random graphs, which has been normalized so that the lower bound is 00 and the upper bound is $1$. For each number of blue robots, and for the blue security strategy and naive strategy, the value is plotted against the number of nodes in the random graphs. The plus sign indicates the mean, the center horizontal line indicates the median value, the middle $50\%$ of the values are inside the box, the whiskers indicate the furthest non-outlier values, and the circles indicate the outliers. Some outliers beyond the axis bounds have their values shown in annotations. The blue player’s NE policy outperforms the baseline security strategy, which in turn outperforms the naive strategy, which cannot guarantee the upper bound on the cost. The values are always between the upper and lower bounds for the NE policies and the security policy, the mean decreases as the number of nodes increases, and the mean decreases with the number of robots.
@@
Figure 7 also plots the values per robot $\tilde{V}_{M}(S(0))$ for the random graphs for different numbers of robots. The mean value per robot decreases as number of robots increases, for each graph size, demonstrating that having more robots is an advantage for the blue player under NE policies.
@@
Figure 8 shows the effect of the red player’s ammo count on the normalized game value. The average of the normalized values is taken over all the random graphs and plotted against the ammo count. The value is nondecreasing with respect to the ammo count, and the value plateaus around 4 ammo, below the upper bound for the value, suggesting that having more ammo stops benefiting the red player after a point.
@@
Figure 8: This shows the average game value, over about $500$ random graphs as a function of ammo, normalized so that the lower bound is 00 and the upper bound is $1$. The value is nondecreasing in ammo and it asymptotically approaches a value which is significantly lower than the upper bound.
@@ VI Conclusions
To handle uncertainty for robots moving through hazardous, time-varying environments in the presence of an adversary, we have formulated a novel zero-sum game where the blue player’s robots move through a graph with time-varying costs to reach a goal, while the red player controls those costs. This can be appropriate for military or security applications, or to model the worst-case natural changes. Due to the complexity of the game, we provided security strategies for both players and a theoretical guarantee that the blue robots reach the goal. We provided methods to simplify computation, and demonstrated the game solution through numerical methods. Both players benefit from mixed strategies, while the blue player benefits from having multiple robots which use coordinated strategies, such as splitting up and waiting to traverse risky edges synchronously, ensuring at least some take the cheaper paths. While this work made the strong assumption of perfect state information for the blue player, future work can include formulating and solving games of incomplete information, where the blue player is uncertain of the red player’s capabilities, and imperfect information, where the blue robots are only aware of changes in the graph close to them, to model the more realistic case of limited perception for the robots.